Should you accept a project that costs $125,000 initially, and provides cash flows of $45,000, $60,000 and $75,000 for the following three years, given that the project has the same risk as another project that requires an initial investment of $50,000 and requires an annual cash flow of $9,000 forever to financially break-even?
Answer #1

Rate of return (risk) = r

Equating PV for second project:

50000 = 9000/(1 + r) + 9000/(1 + r)2 +.....

50000 = 9000/r

r = 18%

Calculating NPV of first project using r = 18%:

NPV = -125000 + 45000/(1 + 18%) + 60000/1.182 + 75000/1.183

NPV = $ 1874

The project should be accepted as NPV > 0